Grade 11Mechanics

If you are to drive 300km to an interview. The interview is at 11:15am then you plan to drive at 100km/h, so you leave at 8:00am to allow some extra time. You drive at the speed for the first 100km, but then construction work forces you to slow to 40km/h for 40km. What would be the least speed needed for the rest of the the trip so as to arrive in time for the interview.

Grace
6 Years agoGrade 11

1 Answer

Arun
6 Years ago
let the diagram as
 
08:00---------------I I-----------------------11.15
start-----------------construct--------... finish
100 km/h-----------40 km/h---------- V km/h-----------
------ 100 km------- 40 km ---------- 140 km -----------
 
total time needed = 11.35 - 08.00 = 3 h 35 m = 3 ⁷/₁₂ h
the 1st 100 km take 100/100 = 1 h
the 2nd 40 km take 40/40 = 1 h
 
the remaining time is 3 ⁷/₁₂ - 2 = 1 ⁷/₁₂ h
 
minimum speed needed is = 150/(1 ⁷/₁₂) ≈ 94.73 km/h